Hi, Could you help me to find a way to this problem?!Tanks=)
One number is 6 less than a secound number.The product of the numbers is 48 more than the square of the smaller number.Find the two numbers.
I tryed to solve with this way but i couldnt finish it:




Answer by RAY100(1637)

(1) x =y-6,,,,,,x = smaller, y = larger
.
(2) xy = 48 + x^2,,,,,product means multiply
.
subst (1) in (2)
.
y(y-6) = 48 + (y-6)^2
.
y^2 -6y = 48 + y^2 -12y +36
.
6y = 84
.
y = 14
.
.
x=y-6 = 14-6 = 8
.
check in (2),,14 * 8 = 48 + 64 = 112,,,,ok